# WebDevJudge: Evaluating (M)LLMs as Critiques for Web Development Quality

### v1

@@ Abstract
The paradigm of LLM-as-a-judge is emerging as a scalable and efficient alternative to human evaluation, demonstrating strong performance on well-defined tasks. However, its reliability in open-ended tasks with dynamic environments and complex interactions remains unexplored. To bridge the gap, we introduce WebDevJudge, a systematic benchmark for assessing LLM-as-a-judge performance in web development, with support for both non-interactive evaluation based on static observations and continuous interactive evaluation with a dynamic web environment. WebDevJudge comprises human preference labels over paired web implementations, annotated with structured and query-grounded rubrics to ensure high-quality ground truth. Using this benchmark, we comprehensively evaluate various evaluators, including LLMs, MLLMs, and agentic workflows. We systematically investigate the impact of different paradigms and guidance mechanisms. Our experiments reveal a significant gap between LLM judges and human experts. In-depth analysis indicates this gap stems from fundamental model limitations, including failures in recognizing functional equivalence, verifying task feasibility, and mitigating bias. Overall, WebDevJudge presents a significant challenge to LLM-as-a-judge, offering insights to guide future research toward developing more reliable and capable automated evaluators for complicated scenarios. Code and data are available at https://github.com/lcy2723/WebDevJudge.
@@ 1 Introduction
Evaluate, refine, then evaluate again and refine again—large language models (LLMs) have achieved remarkable success across various domains through this iterative cycle (Madaan et al., 2023; Chen et al., 2023; Shafayat et al., 2025; Li et al., 2025a). Conventional evaluation paradigms rely heavily on human assessment (Zhong et al., 2024; Starace et al., 2025), which, while meticulous, presents a critical bottleneck due to its high cost and low scalability. In response, the paradigm of LLM-as-a-judge has emerged as a promising alternative (Zheng et al., 2023; Dubois et al., 2024), offering a scalable and cost-effective solution for crucial development stages like verification (Lightman et al., 2024) and reward modeling (Yuan et al., 2025). With the advent of sophisticated language agents capable of planning, tool use, and collaboration (Yao et al., 2023; Qin et al., 2024a; Liang et al., 2024), the role of LLM-as-a-judge is rapidly expanding beyond basic, well-defined tasks to encompass challenging real-world problems (Zhuge et al., 2025; Bian et al., 2025). This progression is critical, as it paves the way for more capable automated evaluators, creating the possibility for language models to self-evolve in complex, real-world applications.
@@
However, a fundamental question regarding the reliability of LLM-as-a-judge persists. Its effectiveness is well-established for static and basic tasks (Chen et al., 2024; Saha et al., 2025; Gou et al., 2025; Lù et al., 2025), but these successes share a critical commonality: they rely on static assessment of final outcomes. The reliability of LLM-as-a-judge in dynamic, open-ended domains involving complex interaction remains largely unexplored. Such contexts introduce significant challenges: dynamic evaluation requires continuous interaction with and comprehension of a changing environment (Li et al., 2023a; Paglieri et al., 2025), while their open-ended nature necessitates establishing feasible assessment standards (Li et al., 2025b). This gap between the expanding scope of automated judges and the lack of rigorous validation in complex, interactive settings highlights an urgent need for a new meta-evaluation benchmark. Such a benchmark is essential to assess, interpret, and ultimately enhance the reliability of LLM-based judges as we approach increasingly autonomous AI systems.
@@
Consequently, in this work, we introduce WebDevJudge, a meta-evaluation benchmark for assessing LLM-as-a-judge on a representative complex and interactive task using the context of web development. Web development offers an ideal testbed for complex dynamic evaluation, as it inherently has interaction requirements, whose assessment depends not just on static code but on real-time interaction. The task is also intrinsically open-ended, lacking a single absolute answer. To establish a high-quality ground truth, we introduce a structured annotation methodology using query-grounded rubric trees, which decomposes high-level requirements into a verifiable hierarchy of fine-grained criteria. This rigorous protocol combining rubric with human judges yields an inter-annotator agreement of 89.7%, which substantially exceeds the 63% reported for MT-Bench (Zheng et al., 2023), confirming the reliability of the preference labels. Departing from traditional benchmarks focused on static text (Tan et al., 2025), WebDevJudge supports both static and interactive evaluation by offering multifaceted representations of each web implementation, including its source code, screenshot of the rendered webpage, and a fully interactive environment for dynamic assessment, as shown in the right part of Figure 1. Evaluator performance is measured by agreement with expert human preferences on paired web implementations—a methodology widely adopted for nuanced, open-ended tasks where the absolute answer is deficient (Bai et al., 2022; Lambert et al., 2025).
@@
Using WebDevJudge, we conduct a comprehensive evaluation of a wide array of judges, including LLMs, MLLMs, and agentic workflows, under various paradigms and guidance mechanisms. Our experiments reveal that a significant capabilities gap persists between the most advanced models and human experts, with a performance discrepancy exceeding 15%. Notably, different guidance strategies provide only marginal improvements in the pairwise comparison setting, suggesting that preference prediction through comparative assessment represents an internalized capability in models (Hua et al., 2025; Yu et al., 2025). Furthermore, while agentic workflows appear well-suited for interactive task evaluation, they fail to outperform vanilla models due to error accumulation across planning and execution stages (Pan et al., 2025). Through detailed error analysis and case studies, we systematically investigate the failure modes of automated evaluators. As part of this analysis, we construct WebDevJudge-Unit, a diagnostic dataset specifically designed to evaluate feasibility verification capabilities of different types of LLM-as-a-judge. Our investigation reveals two fundamental performance bottlenecks: (1) a persistent inability to recognize functional equivalence between diverse implementations that achieve the same objectives through different approaches or terminologies, such as implementations using the same title element with only variations in text, and (2) systematic weaknesses in feasibility verification, where static assessment suffers from low precision due to static code analysis limitations while interactive agents exhibit low recall stemming from their own operational constraints. These compounding errors ultimately undermine evaluator performance, pointing toward fundamental research directions for developing truly reliable automated evaluators in complex, interactive domains. Our main contributions are as follows:
@@
  • Comprehensive empirical evaluation of (M)LLMs and agentic workflow reveals that current LLM-as-a-judge approaches still fall significantly short of human-level reliability.
@@ 2 Related work
The paradigm of LLM-as-a-Judge leverages powerful large language models to simulate human-like assessment, enabling scalable and cost-efficient evaluation (Gu et al., 2025). This approach has been widely adopted across various domains, including question generation and answering (Bai et al., 2023), data filtering (Li et al., 2024; Xu et al., 2025), and trajectory evaluation (Pan et al., 2024; Xue et al., 2025). Typical implementations employ a strong LLM as an evaluator, which compares or scores candidate responses against predefined criteria. Common comparison methods include pairwise comparison and single-answer grading Zheng et al. (2023). As application scenarios broaden, the conventional LLM-as-a-Judge approach is evolving into Agent-as-a-Judge (Zhuge et al., 2025), wherein LLM-based agents are equipped with tool-using and collaboration capabilities. Gou et al. (2025) propose a judge agent with extractor and verifier to evaluate deep research tasks with rubric trees, while Zhuge et al. (2025) design a modular agentic framework to evaluate agentic systems. Despite its efficiency, this method faces several challenges. Position bias (Wang et al., 2024) and verbosity (Ye et al., 2025) preference may skew results, and the creation of detailed, human-written rubrics is both labor-intensive and difficult to scale (Starace et al., 2025). Additionally, typical assessments may overemphasize final outcomes, limiting their applicability in open-ended interactive tasks (Zhang et al., 2025a). To evaluate the judge capability of LLMs, we introduce a benchmark on open-ended web development scenarios. It is designed to assess a wide range of LLM-based evaluators, with the dual purpose of benchmarking their performance and exposing fundamental shortcomings in existing automated evaluation approaches.
@@
Figure 1: Overview of WebDevJudge. Left: Data Collection with filtering. Center: Annotation with Rubric. Right: Evaluate various evaluators based on different evaluation paradigms.
@@ 3.1 Data Filtering
After applying these filters, we retain 1,713 high-quality instances. Further details regarding the data construction process are provided in Appendix A.
Table 1: Agreement rates with and without the rubric-based annotation. The ‘without rubric’ part shows agreements between: (1) annotators and (2) annotators and the original labels. The ‘with rubric’ part shows inter-annotator agreements under rubrics from different sources.
@@ 3.2 Annotation via Rubric tree
We validate the effectiveness of the rubric tree by manually constructing trees for 50 instances and asking two annotators to evaluate these using the rubric as a reference. As shown in Table 1, rubric-guided annotation significantly improves agreement rates. To scale this process, we employ few-shot LLM generation to automatically produce rubric trees. A third annotator labels the same instances using these generated rubrics, achieving high agreement with human-written rubrics—confirming the utility of LLM-generated rubrics. We then use generated rubric trees to annotate the remaining data via two expert annotators with software engineering backgrounds. During annotation, we perform manual inspections to exclude incompatible and harmful cases. Final inter-annotator agreement reaches 89.7% (with ties) and 94.0% (without ties), demonstrating high consistency.
@@ 3.3 Data Statistics
The final benchmark consists of 654 instances, with preference distributions as follows: 258 for $a$, 258 for $b$, and 138 ties. To characterize the query domain coverage, we conduct a topic analysis based on the original topics webdev-arena-preference-10k dataset (Vichare et al., 2025). Fine-grained topics are grouped into three broad categories: Digital Design, Game and App Development, and Web and Specialized Technologies based on their shared characteristics and application contexts of representative examples. Table 2 and Figure 2 provide detailed category distributions and visualizations.
@@ 4.1 Evaluator Performance on WebDevJudge
To enable structured and domain-grounded assessment, we design a multi-dimensional Likert scale inspired by prior work (Lan et al., 2024; Zhang et al., 2025a) and incorporate principles from international software testing standards (ISO/IEC/IEEE, 2022). Our criteria comprises four dimensions, Functionality, UI Quality, Code Quality, and Interactivity, each with several sub-criteria rated on a 5-point Likert scale (1: lowest, 5: highest). Detailed criteria are provided in Appendix B.
@@
Model/Method Digital Design Game & App Web & Special Average Single Pair Single Pair Single Pair Single Pair Vanilla Non-reasoning Models  GPT-4.1 54.95 65.77 61.33 67.56 60.39 64.73 58.87 66.06  GPT-4o 51.80 64.41 56.00 64.44 60.39 64.25 55.96 64.37  Qwen-2.5-VL-72B-Inst. 46.40 64.86 48.44 62.67 52.17 63.29 48.93 63.61  Gemini-2.5-flash-lite 48.20 62.16 48.89 57.78 47.34 62.80 48.17 60.86 DeepSeek-V3-0324 58.11 63.96 54.22 61.33 59.90 65.70 57.34 63.61 Kimi-K2-Inst. 62.16 64.41 56.00 61.78 52.17 65.22 56.88 63.76 Qwen3-235B-A22B-Inst. 53.60 63.51 59.56 64.00 58.45 61.35 57.19 63.00 Qwen3-30B-A3B-Inst. 54.50 64.41 58.67 60.44 55.07 65.22 56.12 63.30 Qwen2.5-32B-Inst. 48.65 61.26 53.33 64.89 51.69 61.35 51.22 62.54 Gemma-3-27B-it 44.14 59.01 52.00 54.67 45.41 60.39 47.25 57.95 Qwen2.5-14B-Inst. 48.20 56.31 44.00 51.56 38.65 52.17 43.73 53.36 Reasoning Models  Claude-4-Sonnet 50.90 67.12 58.67 68.44 57.00 62.32 55.50 66.06  Claude-3.7-Sonnet 65.77 67.12 59.11 61.33 57.49 67.15 60.86 65.14  Gemini-2.5-pro 51.35 61.26 50.22 67.11 43.48 57.97 48.47 62.23 GLM-4.5 57.66 66.22 52.44 64.44 50.72 60.87 53.67 63.91 DeepSeek-R1-0528 50.00 64.41 56.00 63.11 49.76 61.84 51.99 63.15 Qwen3-30B-A3B-Think. 54.05 59.91 53.33 59.11 52.17 60.39 53.21 59.79 Agentic Intention rubrics 37.39 - 32.00 - 47.34 - 38.69 - Static rubrics 40.09 - 38.67 - 43.48 - 40.67 - Dynamic rubrics 52.70 - 55.56 - 50.72 - 53.06 - Combined rubrics 54.50 - 56.44 - 57.49 - 56.12 - Human Pairwise comparison - 85.78 - 85.02 - 83.78 - 84.82
@@
LLM-as-a-judge falls short of human-level reliability on complex evaluation tasks. The primary finding is that no current model achieves a sufficient level of agreement with expert judgments. The top-performing evaluator, Claude-4-Sonnet under the pairwise paradigm, attains an agreement rate of only 66.06%. This substantial gap underscores the inherent complexity of web development evaluation, which demands a holistic assessment of functionality and aesthetics. We also observe a clear performance ceiling: while smaller models exhibit scaling effects, larger and more capable models show diminishing returns, consistently plateauing below the mid-70% agreement rate.

### v2

@@ Abstract
The paradigm of LLM-as-a-judge is emerging as a scalable and efficient alternative to human evaluation, demonstrating strong performance on well-defined tasks. However, its reliability in open-ended tasks with dynamic environments and complex interactions remains unexplored. To bridge the gap, we introduce WebDevJudge, a systematic benchmark for assessing LLM-as-a-judge performance in web development, with support for both non-interactive evaluation based on static observations and continuous interactive evaluation with a dynamic web environment. WebDevJudge comprises human preference labels over paired web implementations, annotated with structured and query-grounded rubrics to ensure high-quality ground truth. Using this benchmark, we comprehensively evaluate various evaluators, including LLMs, MLLMs, and agentic workflows. We systematically investigate the impact of different paradigms and guidance mechanisms. Our experiments reveal a significant gap between LLM judges and human experts. In-depth analysis indicates this gap stems from fundamental model limitations, including failures in recognizing functional equivalence, verifying task feasibility, and mitigating bias. Overall, WebDevJudge presents a challenge to LLM-as-a-judge, offering insights to guide future research toward developing more reliable and capable automated evaluators for complicated scenarios. Code and data are available at https://github.com/lcy2723/WebDevJudge.
@@ 1 Introduction
Evaluate, refine, then evaluate again and refine again—large language models (LLMs) have achieved remarkable success across various domains through this iterative cycle (Madaan et al., 2023; Chen et al., 2023; Shafayat et al., 2025). Conventional evaluation paradigms rely heavily on human assessment (Zhong et al., 2024; Starace et al., 2025), which, while meticulous, presents a critical bottleneck due to its high cost and low scalability. In response, the paradigm of LLM-as-a-judge has emerged as a promising alternative (Zheng et al., 2023; Dubois et al., 2024), offering a scalable and cost-effective solution for crucial development stages like verification (Lightman et al., 2024) and reward modeling (Yuan et al., 2025). With the advent of sophisticated language agents capable of planning, tool use, and collaboration (Yao et al., 2023; Qin et al., 2024a; Liang et al., 2024), the role of LLM-as-a-judge is rapidly expanding beyond basic, well-defined tasks to encompass challenging real-world problems (Zhuge et al., 2025; Bian et al., 2025). This progression is critical, as it paves the way for more capable automated evaluators, creating the possibility for language models to self-evolve in complex, real-world applications.
@@
However, a fundamental question regarding the reliability of LLM-as-a-judge persists. Its effectiveness is well-established for static and basic tasks (Chen et al., 2024; Saha et al., 2025; Gou et al., 2025; Lù et al., 2025), but these successes share a critical commonality: they rely on static assessment of final outcomes. The reliability of LLM-as-a-judge in dynamic, open-ended domains involving complex interaction remains largely unexplored. Such contexts introduce significant challenges: dynamic evaluation requires continuous interaction with and comprehension of a changing environment (Li et al., 2023a; Paglieri et al., 2025), while their open-ended nature necessitates establishing feasible assessment standards (Li et al., 2025). This gap between the expanding scope of automated judges and the lack of rigorous validation in complex, interactive settings highlights an urgent need for a new meta-evaluation benchmark. Such a benchmark is essential to assess, interpret, and ultimately enhance the reliability of LLM-based judges as we approach increasingly autonomous AI systems.
@@
Consequently, in this work, we introduce WebDevJudge, a meta-evaluation benchmark for assessing LLM-as-a-judge on a representative complex and interactive task using the context of web development. Web development offers an ideal testbed for complex dynamic evaluation, as it inherently has interaction requirements, whose assessment depends not just on static code but on real-time interaction. The task is also intrinsically open-ended, lacking a single absolute answer. To establish a high-quality ground truth, we introduce a structured annotation methodology using query-grounded rubric trees, which decomposes high-level requirements into a verifiable hierarchy of fine-grained criteria. This rigorous protocol combining rubric with human judges yields an inter-annotator agreement over 80%, which substantially exceeds the 63% reported for MT-Bench (Zheng et al., 2023), confirming the reliability of the preference labels. Departing from traditional benchmarks focused on static text (Tan et al., 2025), WebDevJudge supports both static and interactive evaluation by offering multifaceted representations of each web implementation, including its source code, screenshot of the rendered webpage, and a fully interactive environment for dynamic assessment, as shown in the right part of Figure 1. Evaluator performance is measured by agreement with expert human preferences on paired web implementations—a methodology widely adopted for nuanced, open-ended tasks where the absolute answer is deficient (Bai et al., 2022; Lambert et al., 2025).
@@
Using WebDevJudge, we conduct a comprehensive evaluation of a wide array of judges, including LLMs, MLLMs, and agentic workflows, under various paradigms and guidance mechanisms. Our experiments reveal that a significant capabilities gap persists between the advanced models and human experts, with a performance discrepancy of about 15%. Notably, different guidance strategies provide only marginal improvements in the pairwise comparison setting, suggesting that preference prediction through comparative assessment represents an internalized capability in models (Hua et al., 2025; Yu et al., 2025). Furthermore, while agentic workflows appear well-suited for interactive task evaluation, they fail to outperform vanilla models due to error accumulation across planning and execution stages (Pan et al., 2025). Through detailed error analysis and case studies, we systematically investigate the failure modes of automated evaluators. As part of this analysis, we construct WebDevJudge-Unit, a diagnostic dataset specifically designed to evaluate feasibility verification capabilities of different types of LLM-as-a-judge. Our investigation reveals two fundamental performance bottlenecks: (1) a persistent inability to recognize functional equivalence between diverse implementations that achieve the same objectives through different approaches or terminologies, such as implementations using the same title element with only variations in text, and (2) systematic weaknesses in feasibility verification, where static assessment suffers from low precision due to static code analysis limitations while interactive agents exhibit low recall stemming from their own operational constraints. These compounding errors ultimately undermine evaluator performance, pointing toward fundamental research directions for developing truly reliable automated evaluators in complex, interactive domains. Our main contributions are as follows:
@@
  • Comprehensive empirical evaluation of (M)LLMs and agentic workflow reveals that current LLM-as-a-judge still fall short of human-level reliability in preference prediction.
@@ 2 Related work
The paradigm of LLM-as-a-Judge leverages powerful large language models to simulate human-like assessment, enabling scalable and cost-efficient evaluation (Gu et al., 2025). This approach has been widely adopted across various domains, including question answering (Bai et al., 2023), data filtering (Li et al., 2024; Xu et al., 2025), and trajectory evaluation (Pan et al., 2024; Xue et al., 2025). Typical implementations employ a strong LLM as an evaluator, which compares or scores candidate responses against predefined criteria. Common comparison methods include pairwise comparison and single-answer grading Zheng et al. (2023). As application scenarios broaden, the conventional LLM-as-a-Judge approach is evolving into Agent-as-a-Judge (Zhuge et al., 2025), wherein LLM-based agents are equipped with tool-using and collaboration capabilities. Gou et al. (2025) propose a judge agent with extractor and verifier to evaluate deep research tasks with rubric trees, while Zhuge et al. (2025) design a modular agentic framework to evaluate agentic systems. Despite its efficiency, this method faces several challenges. Position bias (Wang et al., 2024) and verbosity (Ye et al., 2025) preference may skew results, and the creation of detailed, human-written rubrics is both labor-intensive and difficult to scale (Starace et al., 2025). Additionally, typical assessments may overemphasize final outcomes, limiting their applicability in open-ended interactive tasks (Zhang et al., 2025a). To evaluate the judge capability of LLMs, we introduce a benchmark on open-ended web development scenarios. It is designed to assess a wide range of LLM-based evaluators, with the dual purpose of benchmarking their performance and exposing fundamental shortcomings in existing automated evaluation approaches.
@@
Figure 1: Overview of WebDevJudge. Left: Data Collection with query-based and environment-based filtering. Center: Preference label annotation with verifiable rubric tree. Right: Evaluate (M)LLM-based and agentic evaluators under pairwise and single-answer paradigms.
@@ 3.1 Data Filtering
After applying these filters, we retain 1,713 high-quality instances, we sample 700 instances for further annotation. Details regarding the data construction process are provided in Appendix B.
Table 1: Annotation agreement rates with and without the verifiable rubric. The ‘without rubric’ part shows agreements between: (1) annotators and (2) annotators and the original labels. The ‘with rubric’ part shows inter-annotator agreements under human-written and LLM-generated rubrics.
@@ 3.2 Annotation via Rubric tree
We validate the effectiveness of the rubric tree by manually constructing trees for 50 instances and asking two annotators to evaluate these using the rubric as a reference. As shown in Table 1, rubric-guided annotation significantly improves agreement rates. To scale this process, we employ few-shot LLM generation to automatically produce rubric trees. A third annotator labels the same instances using these generated rubrics, achieving high agreement with human-written rubrics—confirming the utility of LLM-generated rubrics. We then use generated rubric trees to annotate the remaining data via two expert annotators with software engineering backgrounds. During annotation, we perform manual inspections to exclude incompatible and harmful cases. Final inter-annotator agreement on a sampled subset reaches 89.7%, demonstrating high consistency.
@@ 3.3 Data Statistics
The final benchmark consists of 654 instances, with preference distributions as follows: 269 for $a$, 276 for $b$, and 109 ties. To characterize the query domain coverage, we conduct a topic analysis based on the original topics webdev-arena-preference-10k dataset (Vichare et al., 2025). Fine-grained topics are grouped into three broad categories: Digital Design, Game and App Development, and Web and Specialized Technologies based on their shared characteristics and application contexts of representative examples. Table 2 and Figure 2 provide detailed category distributions and visualizations.
@@ 4.1 Evaluator Performance on WebDevJudge
To enable structured and domain-grounded assessment, we design a multi-dimensional Likert scale inspired by prior work (Lan et al., 2024; Zhang et al., 2025a) and incorporate principles from international software testing standards (ISO/IEC/IEEE, 2022). Our criteria comprises four dimensions, Functionality, UI Quality, Code Quality, and Interactivity, each with several sub-criteria rated on a 5-point Likert scale (1: lowest, 5: highest). Detailed criteria are provided in Appendix C.
Model/Method Digital Design Game & App Web & Special Average Single Pair Single Pair Single Pair Single Pair Vanilla Non-reasoning Models  GPT-4.1 57.66 70.72 64.89 73.33 59.90 66.67 60.86 70.34  GPT-4o 53.15 68.02 59.56 70.67 60.39 64.25 57.65 67.74  Qwen-2.5-VL-72B-Inst. 46.85 68.02 51.11 66.22 52.17 64.73 50.00 66.36  Gemini-2.5-flash-lite 50.45 63.51 51.56 60.89 48.31 63.77 50.15 62.69 DeepSeek-V3-0324 60.36 67.12 59.11 66.67 60.39 67.63 59.94 67.13 Kimi-K2-Inst. 62.61 68.02 62.22 67.56 54.59 66.18 59.94 67.28 Qwen3-235B-A22B-Inst. 55.86 66.22 63.11 68.89 60.87 62.80 59.94 66.06 Qwen3-30B-A3B-Inst. 58.56 68.02 63.56 65.33 57.00 65.70 59.79 66.36 Qwen2.5-32B-Inst. 50.45 63.51 54.22 67.56 52.17 63.29 52.29 64.83 Gemma-3-27B-it 45.95 61.71 54.22 57.78 46.38 61.35 48.93 60.24 Qwen2.5-14B-Inst. 48.65 58.11 46.22 54.67 39.61 51.21 44.95 54.74 Reasoning Models  Claude-4-Sonnet 54.05 72.52 63.11 72.00 60.39 65.70 59.17 70.18  Claude-3.7-Sonnet 67.57 70.72 64.44 66.22 59.42 70.05 63.91 68.96  Gemini-2.5-pro 51.80 64.41 53.33 72.00 45.41 60.39 50.31 65.75 GLM-4.5 58.56 71.17 56.89 70.67 51.69 63.77 55.81 68.65 DeepSeek-R1-0528 51.80 69.37 61.33 68.89 50.24 62.32 54.59 66.97 Qwen3-30B-A3B-Think. 53.15 64.86 55.56 64.00 53.14 62.80 53.98 63.91 Agentic Intention rubrics 38.74 - 32.00 - 48.31 - 39.45 - Static rubrics 52.25 - 48.00 - 53.14 - 51.07 - Dynamic rubrics 54.05 - 59.11 - 53.62 - 55.66 - Combined rubrics 59.91 - 58.22 - 63.77 - 60.55 - Human Pairwise comparison - 84.23 - 83.11 - 86.47 - 84.56
@@
LLM-as-a-judge falls short of human-level reliability on complex evaluation tasks. The primary finding is that no current model achieves a sufficient level of agreement with expert judgments. The top-performing evaluator, GPT-4.1 under the pairwise paradigm, attains an agreement rate of 70.34%. This gap underscores the inherent complexity of web development evaluation, which demands a holistic assessment of functionality and aesthetics, especially for the “tie” cases. We also observe a clear performance ceiling: while smaller models exhibit scaling effects, larger and more capable models show diminishing returns, consistently plateauing below the 70% agreement rate.